I've never taken statistics before and am really not sure how to go about even working through these problems and could use some help please and thank you.
In 1997 the U.S. population was approximately 266,574,000 people. A. At the rate of 444 prisoners per 100,000population, about how many prisoners were in the system? B. The projected U.S. population for the year 2020 is 323,724,000. If the rate of prisoners per 100,000 stays the same as in 1997 how many prisoners will be in the systems in 2020?
Thank you for your help.

Answer by stanbon(75887)

In 1997 the U.S. population was approximately 266,574,000 people.
A. At the rate of 444 prisoners per 100,000 population, about how many prisoners were in the system?
-----
# of prisoners = 444[266,574,000/100,000] = 444*2,665.74 = 1,183,589
B. The projected U.S. population for the year 2020 is 323,724,000.
If the rate of prisoners per 100,000 stays the same as in 1997 how many prisoners will be in the systems in 2020?
# of prisoners = population/100,000]*444 = 1,437,335
==============
Cheers,
Stan H.
==============